Find the length for side a in the right triangle, if b=4 and c=5.
Answer by checkley77(12844)

a^2+b^2=c^2
a^2+4^2=5^2
a^2+16=25
a^2=25-16
a^2=9
a=sqrt9
a=3 ans. for side a.